# Story Generation from Visual Inputs: Techniques, Related Tasks, and Challenges

### v1

@@ 3.8. Discussion
These metrics guide the development and refinement of visual story generation models by offering insights into their linguistic fluency, relevance to visual content, and overall narrative coherence. They focus on specific linguistic and content-related aspects, providing quantitative measures for precision, recall, and consensus. They are, however, limited by their focus on surface-level textual similarities rather than deep narrative alignment: since a story can be retold in various ways, employing different perspectives while preserving the original message, the corresponding generated text might deviate significantly from the reference text in terms of specific wording.
@@
Huang et al. (Huang et al., 2016) conducted an analysis to evaluate how well BLEU and METEOR correlate with human judgment in the context of VSG according to different metrics. The results in Table 2 show that METEOR, which places a strong emphasis on paraphrasing, has higher correlation with human judgment than BLEU. Nevertheless, in spite of METEOR showing the highest correlation with human judgment, this value is still low, suggesting that a large gap still exists between automatic metrics and human judgment.
@@
Table 2. Correlations of automatic scores against human judgements, with p-values in parentheses (Huang et al., 2016).
@@
The proliferation of diverse LLMs, such as the GPT (OpenAI, 2024c, b, a), Llama (Touvron et al., 2023; Meta, 2024), Mistral (team, 2023), and Claude (Anthropic, 2024b, a, 2024) families, has led to the development of systematic scoring systems, aiming for objective assessments across a wide array of language understanding tasks. These scoring systems include Multi-turn Benchmark (MT-Bench) (Zheng et al., 2023) and Chatbot Arena (Zheng et al., 2023) briefly described below. While they were initially developed for the evaluation of chatbots by human evaluators, they have been adapted to operate using a LLM as the evaluator achieving results close to those obtained by human evaluators.
@@
As LLMs continue to improve, they show potential in replacing human annotators in many tasks (Gilardi et al., 2023). Metrics to evaluate stories could be developed based on LLMs (Zheng et al., 2023), possibly following the underlying principles discussed above.
@@
MT-Bench is a benchmark tailored to test the multi-turn conversation and instruction-following capabilities of LLM-based chat assistants. It includes 80 high-quality multi-turn questions distributed across diverse categories such as writing, roleplay, reasoning, and knowledge domains. In the MT-Bench evaluation process, the user is presented with two distinct conversations, each generated by a different LLM-based chat assistant and is then tasked with deciding which assistant, A, B, or indicating a tie, better followed the instructions and answered the questions.
@@
Chatbot Arena introduces a crowd-sourcing approach, fostering anonymous battles between LLM-based chatbot models. Users interact with two anonymous models simultaneously, posing identical questions to both and subsequently voting for the model that provides the preferred response. The Elo rating system (Zheng et al., 2023) is employed to rank the models based on their performance in these battles. This rating system assigns a numerical score to each model based on their battle outcomes, adjusting these scores higher or lower after each encounter to reflect their relative ability to satisfy user queries effectively compared to their opponents.
@@
In summary, while current metrics provide insights into specific dimensions of VSG, the development of new metrics may be necessary to holistically evaluate creativity, coherence, and emotional engagement, ensuring a more comprehensive understanding of the capabilities and limitations of these models. We anticipate that LLMs can play a role in the development of such metrics.
4. Story Generation
The use of Artificial Intelligence (AI) for automated story generation has been explored since the 1960s and 1970s. Early approaches include rule-based systems (Section 4.1), plan-based methods (Section 4.3), grammar-based systems (Section 4.2), and case-based reasoning systems (Section 4.4). In recent times, with the availability of large datasets and computational resources the these methods have been replaced by deep learning models: Recurrent Neural Networks (RNNs) (Section 4.5) which have been widely explored for story generation tasks; and LLMs (Section 4.6) which currently achieve state-of-the-art performance in various natural language generation tasks.
4.1. Rule-based Methods
Rule-based methods in narrative generation use a predefined set of rules to create stories. These rules govern the relationships between narrative elements, ensuring consistency within the narrative, aligning with the desired outcomes and narrative styles or genres. However, it is essential to acknowledge the potential limitations in flexibility and creativity, as the story’s possibilities are bounded by the predefined rules and the system’s ability to simulate complex interpersonal dynamics and character development.
TALE-SPIN (1977) (Meehan, 1977) was an influential rule-based method that initiated its storytelling process by presenting a central problem to the main character, derived from foundational states. As the narrative unfolds, the character may encounter various sub-problems and unintended side effects, which also require resolution. The complexity of these challenges can be adjusted by the user, influencing the story’s length and intricacy. This capability to modulate narrative complexity allows for a range of storytelling from straightforward tales to more complex narratives, depending on the inter-character dynamics established by the user. A pivotal aspect of narrative construction of TALE-SPIN is its nuanced handling of character relationships which are essential for the progression and richness of the story. These relationships are categorized into various dimensions such as competition, dominance, familiarity, affection, trust, deceit, and indebtedness. These relationships not only set preconditions for the actions of the characters but also emerge as outcomes of their interactions within the narrative. For instance, the act of one character giving a gift to another can lead to the recipient feeling a sense of affection towards the giver, thereby influencing subsequent narrative developments. Fig. 1 shows an example of a story generated by TALE-SPIN.
[…]
Figure 1. An example story by TALE-SPIN (Meehan, 1977).
IDtension (2003) (Szilas, 2003) merges dramatic storytelling with interactive fiction. This architecture is designed to balance narrative consistency with the creative spontaneity essential for immersive interactive experiences. IDtension is composed of five foundational modules: the world of the story, narrative logic, narrative sequencer, user model, and the theater. Each module plays a role in crafting a narrative that responds and evolves based on user interactions, maintaining an equilibrium between narrative depth and interactive freedom. The world of the story module serves as the narrative’s backbone, housing the essential entities and their states, thereby setting the stage for the unfolding drama. The narrative logic then calculates all possible character actions derived from the narrative ecosystem. The narrative sequencer filters through actions to highlight those with the greatest narrative value, ensuring that each user decision is met with consequences that propel the story forward. The user model captures the user’s current narrative context, guiding the sequencer in tailoring the narrative flow to enhance user engagement. The theater is responsible for displaying the action(s), and manages the interaction between the computer and the user. Fig. 2 shows an example of a story generated by IDtension.
[…]
Figure 2. An example story by IDtension (Szilas, 2003) showing the telegraphic nature of the text.
4.2. Grammar-based Methods
Grammar-based methods in narrative generation, a specialized subset of rule-based approaches, employ formal grammars, such as context-free grammars, to systematically construct stories. The vocabulary elements can be narrative elements like the ones described in the previous section. In contrast with rule-based methods, which emphasize logical consistency and specific narrative outcomes through predefined rules guiding character and plot interactions, these grammars consist of a set of production rules that define how various elements of a story can be combined to form coherent narratives. This method emphasizes linguistic correctness and structural coherence, making it suitable for applications where the narrative needs to adhere to specific stylistic or syntactic standards. However, the creativity and variety of the narratives generated are often limited by the constraints of the grammar used. Grammar-based systems are particularly effective in scenarios where the narrative needs to follow a well-defined and consistent structure.
Different grammars have been proposed that capture narrative patterns such as the three-act structure (Aristotle et al., 1997), Propp’s morphology (Propp et al., 2010), Freytag’s pyramid (Freytag and MacEwan, 1894), the hero’s journey monomyth (Campbell, 2008), among others. Grammars can be designed for specific genres or domains.
Examples of grammar-based methods include systems developed by Grimes (Ryan, 2017) and Gervás (Gervás, 2013). Grimes’ system was one of the pioneering approaches in the 1960s to harness computational methods for automatic story generation. The system would randomly select elements from Propp’s list (Propp, 1968) and then order them properly to generate a coherent story. This was a significant achievement, considering the computational limitations of the time. This system employed a grammar-based technique, drawing inspiration from Propp’s morphology, to craft structured narratives. Gervás (2013), decades after Grimes’ system, computationally modeled Russian folktales. His approach involved representing Propp’s morphology by analyzing character roles and narrative functions. This was implemented as a generative grammar that arranges plot elements according to a chronological sequence of narrative functions.
4.3. Plan-based Methods
Plan-based methods for narrative generation involve using planning algorithms to create a storyline. In this approach, the system formulates a narrative goal and then plans a sequence of events or actions that characters can take to achieve that goal. This method allows for dynamic storytelling, as the system can adapt to different narrative goals and constraints. However, incoherent stories can be generated if the planning algorithm is not sophisticated enough. Plan-based methods are useful for creating stories that require logical progression and can adapt to changing story requirements. Examples of plan-based methods include MEXICA (Pérez y Pérez and Sharples, 2001) and Mimesis (Young et al., 2004).
4.3.1. MEXICA
MEXICA (1999) (Pérez y Pérez and Sharples, 2001) used creative storytelling to generate Aztec mythology tales. Character-based planning drove narrative progression based on changes in character emotions extracted from a case library.
MEXICA operates through two main processes. Initially, it builds knowledge from a collection of predefined stories, using this information to create schemas in its Long-Term Memory (LTM) that encapsulate content and rhetorical knowledge specific to a genre or form of writing. Users can add or modify stories in the system’s archive using a specialized language designed for defining these stories (PSDL) and the actions within them (SADL).
The MEXICA storytelling core is an engagement-reflection cycle. In the engagement phase, MEXICA generates stories by planning a sequence of actions, each with associated post-conditions that alter the story world (e.g., emotional links between characters or changes in their physical positions). These post-conditions are not explicitly stated in the story but guide the narrative’s direction. For instance, the action “princess cures jaguar knight’s injuries” not only progresses the plot but also establishes an emotional link (gratitude) between the characters, influencing future actions.
Characters in MEXICA possess individual “story-world contexts”, which are dynamic records of the post-conditions affecting them. These contexts ensure that each character’s storyline reflects their unique experiences and interactions in the narrative, allowing for complex and divergent storylines based on perspectives and relationships of the characters.
When MEXICA reaches a narrative impasse, that is, when it cannot proceed based on the current story-world context, it enters the reflection phase. Here, it analyzes previous stories for inspiration to overcome the block, assessing novelty and interestingness by comparing the current story’s content and dramatic tension with those of past narratives. This process allows MEXICA to break impasses and continue with story generation, thereby simulating a human writer’s strategy of seeking external inspiration to resolve creative obstacles. Fig. 3 shows a story generated by MEXICA. This example shows how limited pronoun variation and repetitive use of character names cause the narrative to be less fluid.
[…]
Figure 3. An example story by MEXICA (Pérez y Pérez and Sharples, 2001).
4.3.2. Mimesis
Mimesis (2004) (Young et al., 2004) addresses a fundamental challenge in game development: creating compelling and context-sensitive behaviors for characters and objects. Its architecture employs a structured approach to generate and execute narrative plans within interactive game environments. The process begins when the world planner receives a request from the game engine, detailing a narrative problem that requires resolution, such as achieving specified narrative objectives using a predetermined set of actions. In response, the world planner formulates a plan, a structured dataset that sequences actions for characters and other game entities to fulfill the narrative goals outlined in the request. Following the creation of the story world plan, the discourse planner manages non-character elements like camera movements and background music, enhancing the narrative unfolding within the game. The discourse planner then uses the world plan to craft a narrative plan that encapsulates all system and user activities planned in response to the game engine’s initial request. The final step involves the execution manager, which receives the narrative plan and builds a Directed Acyclic Graph (DAG) to map temporal dependencies among the actions. It then navigates through the DAG, initiating ready actions, commanding the game engine to execute them, and monitoring their outcomes to ensure successful completion or to record any failures. We do not include an example story of Mimesis as it is a system designed to generate stories in interactive game environments.
4.4. Case-based Reasoning Systems
Case-based reasoning approaches generate new stories by retrieving and adapting prior story plots stored in a memory of narrative cases. When a new story situation is defined, including elements like setting, characters, and author goals, this situation is matched against the narrative case library to find the most similar prior plots. One or more retrieved plots are adapted to the new situation by modifying characters, events and descriptions, creating a new plot that becomes the framework for the new story (it can also be added to the library for future reuse). This approach allows generating novel narratives by reusing and building upon successful past story structures.
Case-based narrative systems must define how to represent, index, match, adapt, and learn from story cases. Large case libraries, similarity metrics, and adaptation rules enable generating a wide range of creative plot variations.
Examples of case-based reasoning systems include MINSTREL (1993) (Turner, 1993) and Story Plot Generation (2004) (Gervás et al., 2004). The latter uses existing stories and case-based reasoning to create new stories based on user queries. MINSTREL is a creative storytelling system that generated original stories about King Arthur. It combines case-based goal planning with models of drama and emotional responses. Fig. 4 shows an example of a story created by the system.
[…]
It was the spring of 1089, and a knight named Lancelot returned to Camelot from elsewhere. Lancelot was hot tempered. Once, Lancelot lost a joust. Because he was hot tempered, Lancelot wanted to destroy his sword. Lancelot struck his sword. His sword was destroyed.
One day, a lady of the court named Andrea wanted to have some berries. Andrea went to the woods. Andrea had some berries because Andrea picked some berries. Lancelot’s horse moved Lancelot to the woods. This unexpectedly caused him to be near Andrea. Because Lancelot was near Andrea, Lancelot saw Andrea. Lancelot loved Andrea.
Some time later, Lancelot’s horse moved Lancelot to the woods unintentionally, again causing him to be near Andrea. Lancelot knew that Andrea kissed with a knight named Frederick because Lancelot saw that Andrea kissed with Frederick. Lancelot believed that Andrea loved Frederick. Lancelot loved Andrea. Because Lancelot loved Andrea, Lancelot wanted to be the love of Andrea. But he could not because Andrea loved Frederick. Lancelot hated Frederick. Because Lancelot was hot tempered, Lancelot wanted to kill Frederick. Lancelot went to Frederick. Lancelot fought with Frederick. Frederick was dead.
Andrea went to Frederick. Andrea told Lancelot that Andrea was siblings with Frederick. Lancelot believed that Andrea was siblings with Frederick. Lancelot wanted to take back that he wanted to kill Frederick, but he could not because Frederick was dead. Lancelot hated himself. Lancelot became a hermit. Frederick was buried in the woods. Andrea became a nun.
[…]
Figure 4. An example story by MINSTREL (Turner, 1993).
MINSTREL narratives, while coherent, lack several elements that would make them more engaging. The story’s phrasing is repetitive, which, along with the lack of descriptive detail, hampers the reader’s ability to fully immerse themselves in the world or connect with the characters. It also misses opportunities to demonstrate the emotions and motivations of the characters through their actions or dialogues, rather than through plain statements.
4.5. Recurrent Neural Networks
Recurrent Neural Network (RNN) are a deep learning models for sequential data processing, with applications ranging from language modeling to image captioning and text generation, that influenced later models. Sequential processing in RNNs involves handling sentences word by word, which, while effective for capturing short-term dependencies, poses challenges in learning long-range dependencies.
RNNs encounter difficulties with long-range dependencies due to issues like vanishing or exploding gradients. This poses a limitation in capturing complex dependencies over long sequences. Architectures such as Long Short-Term Memory network (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Units (GRUs) (Chung et al., 2014) were introduced to address these issues.
Choi et al. (Choi et al., 2016) used RNNs to generate stories by predicting the next sentence vector based on the current input, ensuring the continuity of the narrative flow within a story. Jain et al.(Jain et al., 2017) used GRUs to encode sequences of variable length input descriptions and to decode them as story-like summaries. Fig. 5(a) shows an example input provided to the model, and Fig. 5(b) shows an example story generated by the model. RNNs base their narratives generation decisions on statistical associations derived from the datasets on which they are trained, meaning that narrative elements are learned implicitly from the training data, leading to outputs that can sometimes be unpredictable. In the example we can see this unpredictability in the repetition of the phrase “we also saw some lily pads .”.
[…]
(a) Input captions.
[…]
(b) Generated story.
Figure 5. A story generated by an RNN (Jain et al., 2017).
4.6. Transformers and Large Language Models
Transformers (Vaswani et al., 2017) are a class of deep-learning models that represent a paradigm shift in sequential data processing and are the basis of the Large Language Models (LLMs) revolution. In contrast with RNNs, which process words sequentially in a recurrent way, transformers process the entire sentence simultaneously, eliminating the risk of “forgetting” past information and improving the training process efficiency.
LLMs are composed of multiple transformer layers, which enable them to attend to information at the word, sentence, and paragraph levels. This multi-layered approach allows for the construction of narratives that are not only coherent but also perceived as richly detailed and creatively varied. LLMs also leverage their extensive training data to produce stories across a wide range of styles and genres. However, as with RNNs, LLMs do not explicitly learn narrative elements or the underlying structure of storytelling. Their ability to generate stories originates from the statistical patterns and associations found within their training corpora. Nevertheless, the transformer architecture’s ability to attend to information across different parts and levels of the text enables LLMs to maintain context and coherence over diverse textual structures, contributing to their effectiveness in generating narrative content.
Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) is one of the first LLMs. It was widely adopted and established several state-of-the-art results for downstream tasks, even though it is now considered a small model: the base version has 110 million parameters, while current state-of-the-art models, such as Generative Pre-trained Transformer (GPT)-4 (OpenAI, 2024b) and Mixtral (team, 2023), have hundreds of billions or even trillions of parameters. Current LLMs, such as GPT (OpenAI, 2024b), have demonstrated proficiency in understanding context, generating coherent text, and even simulating human-like conversational abilities.
Some of the key aspects that characterize a LLM and are crucial for their effectiveness in narrative generation are their size (number of parameters), the nature of the training data, and the context window size. Regarding size, larger models tend to perform better, capturing more complex language nuances. However, increasing model size also requires more computational resources for training and inference. Regarding training data, LLMs benefit from diverse and extensive training datasets: a broader range of topics and writing styles in the training data helps the model generalize well to various tasks. The context window size refers to the number of preceding tokens the model considers when generating the next token. This is a critical feature, as it determines how much of the previous text the model can “remember” and use for making predictions about what comes next. A larger context window allows the model to understand longer dependencies in the text, improving coherence and context awareness. This is especially important for tasks that involve lengthy discussions or complex narratives. However, larger context windows come with increased computational costs and this relationship is quadratic (Vaswani et al., 2017), posing significant challenges in terms of memory and processing power, especially for very large models and has been a key focus of recent research efforts (Ding et al., 2023).
Several works using LLMs have been proposed for story generation. Tang et al. (Tang et al., 2022) proposed TextBox 2.0, a library that uses LLMs to generate stories. We do not include an example of a story generated by TextBox 2.0 as the authors did not provide one in their paper. Yang et al. (Yang et al., 2022) proposed Re3, a framework that prompts a general-purpose LLM to build a story plan consisting of a setting, characters, and a plot. This structured plan is generated from a short sentence describing the story, called premise. Story passages are then generated by repeatedly injecting contextual information from both the plan and current story state via prompt. This approach is one of the first approaches to explicitly model the elements character, setting, and plot in the story generation process. Fig. 6 shows the initial setup and outline for an example premise. The resulting story incorporates all the elements from the premise, including character interactions, setting descriptions, and plot developments. We did not include the generated story as it occupies several pages and is available in the original paper (Yang et al., 2022).
[…]
Premise: A woman discovers that her husband has been cheating on her and decides to get revenge.
Setting: The story is set in the present day, in a small town.
Characters:
1. Diane Chambers is a beautiful woman in her early thirties. She has long dark hair and blue eyes. She is married to Mark Chambers and is a stay-at-home mom.
2. Mark Chambers is a successful businessman in his early forties. He is tall and handsome, with brown hair and green eyes. He is married to Diane Chambers and is the father of two young children.
3. Karen Johnson is a beautiful blonde women in her early thirties. She is Mark Chambers’ mistress and is also married with two young children.
Outline:
1. Diane Chambers discovers that her husband has been cheating on her with another woman.
2. Diane Chambers decides to take revenge on her husband by having an affair of her own.
[…]
Figure 6. Re3 initial setup and outline for the shown premise (Yang et al., 2022).
4.7. Discussion
A critical limitation of rule-based, grammar-based, plan-based, and case-based methods is that they struggle with language variation, especially (but not only) with pronoun variation. This results in the frequent and repetitive use of certain structures and constituents, such as character names, leading to less fluid and less engaging narratives. The principle of “Show, Don’t Tell” (Dietrich and Sundell, 1983) is often underused: showing is more powerful than telling because it allows the reader to experience the story directly, rather than being told what to feel. However, these systems tend to rely on direct exposition to convey character emotions and story developments, rather than implying them through actions and interactions. This failure is not because the methods cannot solve these issues. Rather, they would require an impractically large number of conditions and rules to be manually encoded. RNNs are not immune to these challenges, but their inherent adaptability and capacity for learning from extensive datasets make them more capable of capturing the complexities of narrative storytelling. Likewise, LLMs are able to create fluid narratives with varied sentence structures, dialogues between characters, rich and descriptive details about the setting and characters, because they are able to learn aspects of the world from their training data. This allows them to generate details about the setting and characters that are not explicitly provided in the prompt. LLMs, however, are not without their limitations: they may suffer from hallucinations, that is, generating details that completely deviate from the original prompt, or generate contradictory or nonsensical information.
5. From NLP to computer vision
This section provides an overview of areas at the intersection of computer vision and Natural Language Processing (NLP) that are relevant to the topic of VSG, namely image and video captioning and Visual Question Answering (VQA). These tasks share a common challenge: they require the understanding of visual content and the ability to generate text based on that visual content. Some of the techniques used in these tasks have been adapted to VSG. Special attention is given to approaches based on deep learning architectures.
5.1. Image and Video Captioning

### v2

@@ 3 Benchmark Datasets and Evaluation Metrics
The evaluation metrics, used to assess the performance of VSG models, provide quantitative measures of the quality of generated narratives. Most of these metrics were originally developed for machine translation or summarization but have been adopted by the community for other tasks such as image captioning and story generation: instead of comparing a generated text with a reference translation or summary, the generated text is compared with a reference caption in the case of image captioning, or with a reference story in the case of story generation. These metricsa include Bilingual Evaluation Understudy (BLEU) Papineni et al. (2002), Self-BLEU Zhu et al. (2018), Metric for Evaluation of Translation with Explicit ORdering (METEOR) Banerjee and Lavie (2005), Consensus-based Image Description Evaluation (CIDEr) Vedantam et al. (2014), Recall-Oriented Understudy for Gisting Evaluation (ROUGE) Lin (2004), Perplexity Jurafsky and Martin (2023) and Bert Score Zhang et al. (2019). These metrics guide the development and refinement of VSG models by offering insights into their linguistic fluency, textual diversity, and mostly lexical overlap with reference texts. They focus on specific linguistic and content-related aspects, providing quantitative measures for precision, recall, and consensus. However, they are limited by their focus on surface-level textual similarities rather than deep narrative alignment: since a story can be retold in various ways, employing different perspectives while preserving the original message, the corresponding generated text might deviate significantly from the reference text in terms of specific wording.
@@
Huang et al. Huang et al. (2016) conducted an analysis to evaluate how well BLEU and METEOR correlate with human judgment in the context of VSG according to different metrics. The results in Table 2 show that METEOR, which places a strong emphasis on paraphrasing, has higher correlation with human judgment than BLEU. Nevertheless, in spite of METEOR showing the highest correlation with human judgment, this value is still low, suggesting that a large gap still exists between automatic metrics and human judgment.
@@
Table 2: Correlations of automatic scores against human judgements, with p-values in parentheses Huang et al. (2016).
@@
The proliferation of diverse Large Language Models (LLMs), such as the GPT OpenAI (2024, 2024a, 2024b, 2024c), Llama Touvron et al. (2023); Meta (2024), Mistral team (2023); Mistral AI (2024), and Claude Anthropic (2025, 2024a, 2024b, 2024) families, has led to the development of systematic scoring systems, aiming for objective assessments across a wide array of language understanding tasks. These scoring systems include Multi-turn Benchmark (MT-Bench) Zheng et al. (2023) and Chatbot Arena Zheng et al. (2023) briefly described below. While they were initially developed for the evaluation of chatbots by human evaluators, they have been adapted to operate using a LLM as the evaluator achieving results close to those obtained by human evaluators.
@@
As LLMs continue to improve, they show potential in replacing human annotators in many tasks Gilardi et al. (2023). Metrics to evaluate stories could be developed based on LLMs Zheng et al. (2023), possibly following the underlying principles discussed above.
@@
MT-Bench is a benchmark tailored to test the multi-turn conversation and instruction-following capabilities of LLM-based chat assistants. It includes 80 high-quality multi-turn questions distributed across diverse categories such as writing, roleplay, reasoning, and knowledge domains. In the MT-Bench evaluation process, the user is presented with two distinct conversations, each generated by a different LLM-based chat assistant and is then tasked with deciding which assistant, A, B, or indicating a tie, better followed the instructions and answered the questions. For VSG evaluation, this framework could be adapted by using an LLM as an automated judge to compare stories generated by different models, evaluating them based on criteria such as narrative coherence, visual grounding, and creativity.
@@
Chatbot Arena introduces a crowd-sourcing approach, fostering anonymous battles between LLM-based chatbot models. Users interact with two anonymous models simultaneously, posing identical questions to both and subsequently voting for the model that provides the preferred response. The Elo rating system Zheng et al. (2023) is employed to rank the models based on their performance in these battles. This rating system assigns a numerical score to each model based on their battle outcomes, adjusting these scores higher or lower after each encounter to reflect their relative ability to satisfy user queries effectively compared to their opponents. A similar automated evaluation system for VSG could employ an LLM judge to assess pairs of generated stories given the same image sequence, ranking models based on story quality across multiple dimensions without requiring human annotators.
@@
In summary, while current metrics provide insights into specific dimensions of VSG, the development of new metrics may be necessary to holistically evaluate creativity, coherence, and emotional engagement, ensuring a more comprehensive understanding of the capabilities and limitations of these models. We anticipate that LLMs can play a role in the development of such metrics, serving as automated judges that evaluate stories based on multiple criteria.
4 From NLP to computer vision
This section provides an overview of areas at the intersection of computer vision and Natural Language Processing (NLP) that are relevant to the topic of VSG, namely image and video captioning and Visual Question Answering (VQA). These tasks share a common challenge: they require the understanding of visual content and the ability to generate text based on that visual content. Some of the techniques used in these tasks have been adapted to VSG.
4.1 Image and Video Captioning